# How Do Multilingual Models Remember? Investigating Multilingual Factual Recall Mechanisms

### v1

@@ Appendix E Patching
Table 3: Proportion of times an object is predicted in the other language. In parenthesis the number of examples corresponding to the percentage. In bold when $o_{s}^{l_{t}}$ or $o_{t}^{l_{s}}$ are detected more often for each of the experiments. The total number of examples varies, see total numbers in Table 2.
E.1 Different Relation, Different Subject
@@
Figure 20: Probability of the patch answer and the context answer when patching at different layers. Left is XGLM, right mT5.
E.2 Same Relation, Different Subject
@@
Figure 21: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers. Left is XGLM, right mT5.
@@
E.3 Different Relation, Same Subject
Figure 26: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Top is XGLM, bottom mT5.
@@ E.3 Different Relation, Same Subject
Figure 31: Number of times the patch object is predicted in the context language for the experiment of same relation different subject. Left XGLM, right mT5.
Figure 32: Number of times the object is predicted in the opposite language in mT5 in the different relation same subject experiment.
Figure 33: Number of times the context object is predicted in the patch language in XGLM in the experiment of different relation same subject.

### v2

@@ Appendix F Patching
Table 5: Proportion of times an object is predicted in the other language in the patching experiments with $\{\neq\mathcal{L},=r,\neq s\}$. In parenthesis the number of examples corresponding to the percentage. In bold when $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are detected more often for each of the experiments. The total number of examples varies, see total numbers in Table 4.
XGLM EuroLLM mT5
Patch - Context All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ All $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$
en-es 324 284 298 413 360 391 263 253 249
en-vi 336 329 333 105 102 100 209 207 204
en-tr 88 77 84 111 102 110 209 207 204
en-ru 218 216 218 420 415 419 191 191 190
en-uk 35 35 35 351 347 350 136 136 133
en-ko 46 46 46 98 98 98 165 165 165
en-ar 212 212 212 391 390 391 159 159 159
en-he - - - 19 19 19 166 166 166
en-ja - - - 1 1 1 85 85 85
en-fa - - - - - - 114 114 114
Table 6: Total number of patch-context examples considered in the patching experiments with $\{\neq\mathcal{L},\neq r,=s\}$. The $\mathcal{L}_{c}(o_{p})$ column is the total number of examples where the detection of $\mathcal{L}_{c}(o_{p})$ would be unambiguous, that is, $\mathcal{L}_{c}(o_{p})\neq\mathcal{L}_{p}(o_{p})$, conversely for the $\mathcal{L}_{p}(o_{c})$ column.
XGLM EuroLLM mT5
Patch - Context $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$ $\mathcal{L}_{c}(o_{p})$ $\mathcal{L}_{p}(o_{c})$
en-es 18.0% (51) 3.7% (11) 41.7% (150) 2.8% (11) 0.8% (2) 7.6% (19)
en-vi 9.1% (30) 6.6% (22) 3.9% (4) 8.0% (8) 1.0% (2) 12.3% (25)
en-tr 36.4% (28) 6.0% (5) 56.9% (58) 4.5% (5) 2.9% (6) 9.3% (19)
en-ru 64.8% (140) 4.1% (9) 61.7% (256) 9.3% (39) 6.3% (12) 8.9% (17)
en-uk 17.1% (6) 17.1% (6) 66.0% (229) 2.3% (8) 3.7% (5) 9.0% (12)
en-ko 41.3% (19) 21.7% (10) 71.4% (70) 5.1% (5) 2.4% (4) 12.1% (20)
en-ar 40.1% (85) 1.9% (4) 52.8% (206) 0.8% (3) 1.9% (3) 7.5% (12)
en-he - - 0.0% (0) 0.0% (0) 3.6% (6) 15.7% (26)
en-ja - - 0.0% (0) 0.0% (0) 0.0% (0) 25.9% (22)
en-fa - - - - 2.6% (3) 15.8% (18)
Table 7: Proportion of times an object is predicted in the other language in the patching experiments with $\{\neq\mathcal{L},\neq r,=s\}$. In parenthesis the number of examples corresponding to the percentage. In bold when $\mathcal{L}_{c}(o_{p})$ or $\mathcal{L}_{p}(o_{c})$ are detected more often for each of the experiments. The total number of examples varies, see total numbers in Table 6.
F.1 Different Relation, Different Subject
@@
Figure 32: Probability of the patch answer and the context answer when patching at different layers. Models from top to bottom: XGLM, EuroLLM, mT5.
F.2 Same Relation, Different Subject
@@
Figure 33: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers. Models from top to bottom: XGLM, EuroLLM, mT5.
@@
Figure 36: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers in EuroLLM, for examples with $\{\neq\mathcal{L},=r,\neq s\}$.
Figure 37: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in EuroLLM, for examples with $\{\neq\mathcal{L},=r,\neq s\}$. Note that the plots do not share the y-axis.
@@
F.3 Different Relation, Same Subject
Figure 40: Probability of the patch answer $\mathcal{L}_{p}(o_{p})$ when patching at different layers for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Models from top to bottom: XGLM, EuroLLM, mT5.
@@ F.3 Different Relation, Same Subject
Figure 43: Probability of the $\mathcal{L}_{p}(o_{p})$ when patching at different layers in EuroLLM, , for examples with $\{\neq\mathcal{L},\neq r,=s\}$.
Figure 44: Probability of the $\mathcal{L}_{c}(o_{c})$ when patching at different layers in EuroLLM, for examples with $\{\neq\mathcal{L},\neq r,=s\}$. Note that the plots do not share the y-axis.
@@
Figure 47: Number of times the patch object is predicted in the context language for the experiment of same relation different subject. Models from top to bottom: XGLM, EuroLLM, mT5.
Figure 48: Number of times the object is predicted in the opposite language in mT5 in the $\{\neq r,=s,\neq\mathcal{L}\}$ experiment.
Figure 49: Number of times the context object is predicted in the patch language in the $\{\neq r,=s,\neq\mathcal{L}\}$ experiment. Left XGLM, right EuroLLM.